I need help with the following word problem:
Janine bikes 3 mph faster than her sister, Jessica. Janine can ride
36 miles in 1 hr less time than Jessica can ride the same difference. Find each of their speeds.
Here's what I have so far:
36/x - 1 = 36/x-3
36/x-1x/1x = 36/x-3
(36-x)(x-3) = (36) (x)
36x-108-x^2+3x = 36x
-x^2+3x-108 = 0
I cannot figure out how to factor the -x^2+3x-108.
Please help!

Answer by Alan3354(69443)

Janine bikes 3 mph faster than her sister, Jessica. Janine can ride
36 miles in 1 hr less time than Jessica can ride the same difference. Find each of their speeds.
Here's what I have so far:
36/x - 1 = 36/x-3
36/x-1x/1x = 36/x-3
(36-x)(x-3) = (36) (x)
36x-108-x^2+3x = 36x
-x^2+3x-108 = 0
----------------
Janine can ride
36 miles in 1 hr less time than Jessica can ride the same difference (distance). Find each of their speeds.
x = Jessica's speed
x+3 is Janine's speed
-------
36/(x+3) = (36/x) - 1 = (36-x)/x
36x = (x+3)*(36-x) = -(x+3)*(x-36)
-36x = x^2 - 33x - 108
x^2 + 3x - 108 = 0
(x+12)*(x-9) = 0
x = 9